OpenStudy (anonymous):

amplitude period and phase shift of y=5cos(3x-pi/4)

OpenStudy (anonymous):

Amplitude is 5

OpenStudy (anonymous):

Is there any brackets

OpenStudy (anonymous):

thanks.... yes before 3x and after pi/4

OpenStudy (anonymous):

Then pi/12

OpenStudy (anonymous):

ok is that the period

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok thank you

OpenStudy (aum):

If y = Acos(Bx + C), then A is the Amplitude (2*pi) / B is the Period -C/B is the phase shift. y=5cos(3x-pi/4) Amplitude = 5 Period = (2*pi) / 3 = \(\large \frac 23 \pi\) Phase Shift = -(-pi/4)/3 = \(\large \frac {\pi}{12}\)